Jaime owns a business making decorative boxes to store jewelry, mementos,
and other valuables. The function y = x2 + 50x + 1800 models the profit y that Jaime
has made in month x for the first two years of his business.
a. Write an equation representing the month in which Jaime’s profit is $2400.
b. Use completing the square to find out in which month Jaime’s profit is $2400.

Answer by nerdybill(7384)

Jaime owns a business making decorative boxes to store jewelry, mementos,
and other valuables. The function y = x2 + 50x + 1800 models the profit y that Jaime has made in month x for the first two years of his business.
a. Write an equation representing the month in which Jaime’s profit is $2400.
set y to 2400
2400 = x^2 + 50x + 1800
b. Use completing the square to find out in which month Jaime’s profit is $2400.
2400 = x^2 + 50x + 1800
600 = x^2 + 50x
complete square on right, by adding (b/2)^2 to both sides
600 + 625 = x^2 + 50x + 625
1225 = x^2 + 50x + 625
1225 = (x+25)^2
35 = x+25
10 = x (tenth month)